OpenStudy (anonymous):

.Use the distributive property to rewrite the expression and solve. 8 63 = 8____ + 8 3 = ____

OpenStudy (shaadykhamlichi):

hhhhh again

OpenStudy (anonymous):

yep

OpenStudy (shaadykhamlichi):

8*63=8*60+8*3=504